OpenStudy (anonymous):

find the x in these 45 degrees -45 degrees -90

OpenStudy (e.mccormick):

This is probably taling about the trig ratios where \[\sin\theta=\frac{y}{r}\\ \cos\theta=\frac{x}{r}\\ \tan\theta=\frac{y}{x}\] So you need to find the correct values on the unit circle, then find their x.

OpenStudy (anonymous):

just remember that in any 45 45 90 degree triangle, the length of any side equals one half the hypotenuse times the square root of 2